### v2

# Fairness in AI-Driven Recruitment: Challenges, Metrics, Methods, and Future Directions

Abstract
The recruitment process is crucial to an organization’s ability to position itself for success, from finding qualified and well-fitting job candidates to impacting its output and culture. Therefore, over the past century, human resources experts and industrial-organizational psychologists have established hiring practices such as attracting candidates with job ads, gauging a candidate’s skills with assessments, and using interview questions to assess organizational fit. However, the advent of big data and machine learning has led to a rapid transformation in the traditional recruitment process as many organizations have moved to using artificial intelligence (AI). Given the prevalence of AI-based recruitment, there is growing concern that human biases may carry over to decisions made by these systems, which can amplify the effect through systematic application. Empirical studies have identified prevalent biases in candidate ranking software and chatbot interactions, catalyzing a growing body of research dedicated to AI fairness over the last decade. This paper provides a comprehensive overview of this emerging field by discussing the types of biases encountered in AI-driven recruitment, exploring various fairness metrics and mitigation methods, and examining tools for auditing these systems. We highlight current challenges and outline future directions for developing fair AI recruitment applications, ensuring equitable candidate treatment and enhancing organizational outcomes.
Index Terms:
Artificial intelligence, bias, ethical hiring, fairness, recruitment
I Introduction
The recruitment process is crucial for an organization’s success, influencing both the selection of qualified candidates and the overall organizational culture and performance. Traditional hiring practices, such as job advertisements, skill assessments, and personality tests [1], have been extensively studied by human resource (HR) experts and industrial-organizational (I-O) psychologists. These practices significantly impact organizational outcomes, including performance, employee turnover, and productivity [2, 3]. Furthermore, these outcomes affect millions of job applicants each year, given that the median employee tenure in the United States is 4.2 years and individuals hold an average of 12.3 jobs over their careers [4]. In response to evolving needs, organizations are increasingly adopting innovative recruitment strategies to enhance long-term success and optimize the process of hiring the right jobseekers.
Over the past thirty years, the hiring process has experienced a significant transformation, progressively shifting towards digital recruiting. This evolution can be categorized into three distinct phases. The first phase, digital recruiting 1.0, emerged in the mid-to-late 1990s, marked by the advent of online job descriptions and the ability for jobseekers to submit electronic resumes and applications [5]. The second phase, digital recruiting 2.0, began in the following decade with the rise of job board sites like Indeed and Monster [5]. These platforms aggregated job postings from various organizations, offering jobseekers search and recommendation features to identify positions aligning with their interests and career goals [5]. The current phase, digital recruiting 3.0, has been driven by the integration of artificial intelligence (AI), which has revolutionized recruitment practices across job boards, online networking platforms, and HR applications [5]. Today, AI is extensively utilized by organizations for diverse functions, including job advertisement creation, applicant tracking systems, conducting video and phone interviews, and implementing gamified assessments [6, 1, 5].
I-A Motivation
The growing reliance on AI in recruitment is driven by its demonstrated effectiveness in various HR tasks. A key area is job applicant screening, where digital methods have been shown to expedite the process while potentially reducing human bias [6]. For large organizations handling thousands or even millions of applications, automated recruitment methods have become increasingly essential [5, 7, 6]. AI also plays a significant role in crafting job advertisements, with large language models (LLMs) like ChatGPT being used to draft and outline the necessary skills and qualifications, thereby attracting suitable candidates [8]. Additionally, LLMs are employed to generate interview questions and refine communications with job candidates [8]. Given the significant shift towards remote work, particularly following the COVID-19 pandemic, the reliance on AI for hiring and recruitment decisions is expected to increase substantially [9, 10]. This integration of AI into recruitment processes is underscored by recent statistics showing that, as of 2019, 88% of organizations globally utilized some form of AI in their recruitment activities. Specifically, 41% used AI-based chatbots for candidate engagement, 44% used AI for candidate identification from social media and other public data, and 43% used AI for recommending training to candidates [11]. With the increasing availability of cloud storage and computing resources, the prevalence of AI applications in recruitment is expected to continue growing.
However, as AI adoption becomes more widespread, concerns are growing that decisions made by these systems could be influenced by the biases of organizational personnel or model developers, as evidenced by several recent incidents [1]. For instance, in 2015, Google’s job recommendation system exhibited gender bias by displaying high-income job postings more frequently to men than to women [12]. Similarly, in 2017, Amazon’s AI-based candidate evaluation tool was discontinued due to its discriminatory practice of assigning lower scores to women’s resumes, attributed to the underrepresentation of female applicants in the model’s training data [13, 1]. Moreover, in 2019, Facebook’s housing and job ad delivery system faced similar issues, with job ads being skewed based on users’ gender and race [14]. These examples illustrate how biases commonly found in the hiring process, such as hiring discrimination [15], can easily transfer to AI-based systems through the data used for training algorithms. A timeline of well-known instances of bias in AI related to recruitment applications is shown in Figure 1. Therefore, AI bias in HR applications is already a significant problem, and as AI use continues to grow rapidly, its importance will escalate dramatically in the future.
I-B Related Work and Our Contributions
This paper provides a comprehensive review of fairness in machine learning (ML), deep learning, and AI-based systems specifically for recruitment and hiring, building upon our previous survey [1]. Past studies and surveys have explored biases in non-AI-based recruitment, identifying issues such as confirmation bias (making decisions within the first few minutes of meeting a candidate) and expectation anchor bias (being influenced by a single piece of information), as well as examining how and where AI can be applied in recruitment [16, 17, 18, 19, 20, 21, 22]. Additionally, past research has reviewed specific components of the hiring process, such as AI-driven assessments [23], automated video interviews [24], and their perceptions by job candidates [22].
However, there remains a need for a systematic review of the literature on algorithmic fairness in AI-based recruitment, encompassing all stages of the hiring process where AI can be applied. This need is underscored by the increasing prevalence of AI in recruitment, as evidenced by a 2022 survey revealing that up to 86% of employers reported using technology-driven job interviews, with many adopting automated video interviews as a preferred method [25]. Further examples of AI utilization include using LLMs to craft inclusive job descriptions, employing chatbot systems to engage candidates prior to direct recruiter interaction, and integrating AI into applicant tracking systems for resume screening [8, 6, 1, 5]. Additionally, a 2024 survey indicated that one in four recruiters were already utilizing AI in some capacity [26].
This paper addresses the gap in literature by providing a systematic review of algorithmic fairness in AI-based recruitment, making several key contributions. First, we explore the definitions of fairness, trust, and justice, drawing from I-O psychology, HR practices, and legislative frameworks to provide a foundational understanding for analyzing bias in AI-driven recruitment (Section II). Second, we discuss the ethical and legal principles governing fairness in AI, highlighting current regulations and guidelines that inform fair AI practices in recruitment (Section II-B). Third, we identify and categorize various causes of bias in AI systems, including biases in training data, label definitions, feature selection, proxies, and masking (Section II-C). Our examination extends across all stages of the recruitment process—sourcing, screening, interviewing, and selection—demonstrating how bias can be introduced and mitigated at each step (Section III). Fourth, we review diverse fairness metrics, including individual fairness, group fairness, procedural fairness, and fairness in regression, discussing their relevance for employers, jobseekers, and recruitment platforms (Section IV-A). Fifth, we categorize and evaluate bias mitigation strategies into pre-processing, in-processing, and post-processing methods, highlighting their significance and application in ensuring fair AI-driven recruitment practices (Section IV-B). Sixth, we cover the limitations of existing tools and methods for bias detection and mitigation (Section IV), and outline future challenges in this domain (Section IV-C). Finally, Section V presents concluding remarks and summarizes the key findings. By addressing these critical aspects, our paper aims to set a comprehensive foundation for advancing equitable AI practices in recruitment.
II Background
Before measuring bias, it is essential to define the concept of “fairness” to ensure consistent comparison and evaluation of AI in hiring. The following sections provide definitions of fairness within the context of organizational hiring (Section II-A), outline relevant legislation defining fairness (Section II-B), and discuss common causes of bias in AI systems (Section II-C).
Figure 1: Timeline of reports identifying bias in AI systems related to or influencing current recruitment applications [13, 27, 12, 28, 29, 14, 30, 31, 32, 33, 34, 35, 36, 37].
II-A Relating Fairness to Trust and Justice
The terms “fairness,” “trust,” and “justice” are closely related and often used interchangeably [38]. Before defining fairness, it is useful to examine its relationship with trust and justice, particularly in the context of AI-based hiring.
First, trust is defined in I-O psychology as a psychological state of belief regarding how a system, person, or organization will act in the future based on past interactions [39, 38]. Trust takes time to build and becomes more significant when perceived uncertainty and risk increase (e.g., when the risk of job loss is high) [38]. However, trust can be easily lost, and once lost, it is very difficult to rebuild, leading workers and employees to question the fairness of organizational actions [38]. Therefore, trust in AI-based hiring significantly influences a jobseeker’s perception of the fairness of an organization’s recruitment process.
Trust in technology comprises human characteristics (e.g., personality and ability), environmental characteristics (e.g., cultural background and institutional factors), and technological characteristics (e.g., performance, such as reviews from other users, and goal congruence, ensuring the AI’s goals align with human/organizational goals) [40]. Specific factors that impede trust in AI include its role in technological displacement, lack of transparency, and privacy concerns arising from the collection and reliance on large training datasets [40]. A 2019 poll highlights the lack of trust in AI, revealing that 88% of Americans are skeptical of AI-driven recruitment [41, 42], and as of 2023, 71% opposed AI making final hiring decisions [43]. Therefore, even if model developers consider fairness metrics and precautions, a jobseeker’s trust will significantly affect the perceived fairness of any AI-based decisions.
Organizational justice, in this context, refers to the just treatment in organizational processes and outcomes, such as recruitment [38]. Organizational justice is crucial because it significantly influences perceptions of fairness [38]. The literature identifies three primary types of organizational justice: distributive, procedural, and interactional. Distributive justice pertains to the perceived fairness of the allocation of outcomes or rewards to individuals within an organization. This can be based on the merit or equity norm (i.e., those who work harder or produce more should receive greater rewards), the need norm (i.e., rewards are distributed in proportion to the needs of individuals), or the equality norm (i.e., all individuals receive equal rewards, regardless of effort or need) [38, 44]. Procedural justice is the perceived fairness of the processes by which rewards are distributed [38, 44]. Lastly, interactional justice concerns the perceived extent to which an organization treats an individual or employee with respect. Interactional justice has two aspects: interpersonal justice, which pertains to the extent of respect and politeness accorded to individuals, and informational justice, which reflects how well an organization explains its procedures and outcomes to individuals, ensuring that communication is candid, thorough, and timely [38].
In the recruitment process, these aspects of trust and dimensions of organizational justice all factor into how fairness is defined according to AI fairness research and past legislation, as will be further described.
II-B Fairness According to Legislation and Ethical Principles
In the ML community, the FAT/ML (fairness, accountability, and transparency) group has established five guiding principles to aid developers and product managers in designing and implementing accountable automated decision-making algorithms: responsibility, explainability, accuracy, auditability, and fairness [45]. The group also recommends adherence to OECD privacy principles and the ethical principles outlined in the Belmont Report for human subjects research to ensure algorithmic accountability [46].
Additionally, in 2020, the U.S. Department of Defense defined five ethical principles for its AI systems: responsible, equitable, traceable, reliable, and governable [47, 48]. Following this, numerous companies, including Google [49] and OpenAI [50], have established similar guidelines to mandate the ethical development and integration of their AI technologies.
Similarly, legislation regulating technology has emerged to enhance fairness in AI systems. For instance, in 2018, the European Union enforced the General Data Protection Regulation (GDPR) to protect individuals concerning the privacy, processing, and free movement of personal data [51]. Moreover, in 2021, the European Commission established the first legal framework on AI, categorizing AI systems into different levels of risk [52]. These levels include minimal risk, such as AI-enabled video games or spam filters that pose minimal or no risk to a citizen’s rights or safety; limited risk, where AI systems like chatbots inform users they are interacting with a machine and enable them to make informed decisions before continuing use; high-risk, covering AI systems critical to infrastructure, vocational training, and essential public services like credit scoring, law enforcement, and migration; and unacceptable risk, involving AI systems that present a clear threat to the safety and rights of people [52].
The National Institute of Standards and Technology (NIST) and the Federal Trade Commission (FTC) have also emphasized the need to identify and manage bias in AI [53, 54]. Specifically, the FTC has warned against the sale of racially biased systems that could prevent individuals from obtaining employment, insurance, or other benefits. The FTC highlighted three laws for AI developers and users: Section 5 of the FTC Act, which prohibits unfair or deceptive practices; the Fair Credit Reporting Act, which ensures algorithms do not unfairly deny individuals housing or credit; and the Equal Credit Opportunity Act, which prohibits algorithms from making biased credit decisions based on race, color, religion, national origin, age, sex, marital status, or public assistance status [53, 55].
The Biden Administration launched the National AI Research Resource Task Force [56], as directed by Congress in the National AI Initiative Act of 2020, to advance AI research and explore AI’s implications [57]. Following this, additional legislation such as the Algorithmic Fairness Act of 2020 [58] emerged, and in 2021 alone, 17 states introduced AI regulations [59]. Notably, New York City enacted a law (effective in 2023) prohibiting the use of automated decision-making tools to screen job candidates without prior bias audits and transparency regarding their use [60]. As of 2023, 25 U.S. states have introduced AI bills, and 18 states have adopted or enacted AI legislation [61].
Legislation specific to fairness in the employment and hiring context is enforced by the U.S. Equal Employment Opportunity Commission (EEOC). These laws make it illegal to discriminate against individuals based on race, color, religion, national origin, sex, sexual orientation, or gender identity or expression, as stipulated in Title VII of the Civil Rights Act of 1964, amended by The Civil Rights Act of 1991. Additionally, the Pregnancy Discrimination Act amendment to Title VII protects women against discrimination due to pregnancy or childbirth. The Equal Pay Act of 1963 prohibits wage and benefits discrimination based on sex. The Age Discrimination in Employment Act of 1967 protects individuals 40 years of age or older against age discrimination. The Americans with Disabilities Act of 1990 (Title I) and Sections 501 and 505 of the Rehabilitation Act of 1973 safeguard qualified individuals with disabilities. The Genetic Information Nondiscrimination Act of 2008 prohibits discrimination based on genetic information [62]. Furthermore, the Fourteenth Amendment to the U.S. Constitution provides due process protections to workers [63]. In 2021, the EEOC launched an initiative to ensure that AI and tools used in hiring and employment comply with these laws [64], and as of 2023, it has provided further guidelines for employers [65].
Previously, for employee selection, the EEOC established a requirement to ensure the fair treatment of job candidates by preventing adverse impact due to hiring decisions [66, 67]. According to EEOC guidelines, adverse impact is determined by applying the four-fifths or eighty percent rule. This rule states that the selection rate for a protected group must be at least four-fifths (or 80%) of the rate for the group with the highest selection rate [67]. For instance, as illustrated in Table I, if the selection rate for black applicants is 50% of the selection rate for white applicants, it falls below the 80% threshold, indicating an adverse impact [67].
Applicants Hired Selection Rate Percent Hired
80 White 48 48/80 60%
40 Black 12 12/40 30%
TABLE I: Example selection rate disparity showing adverse impact.
The four-fifths rule has recently come under scrutiny, with proposed legislation advocating for a smaller disparity in selection rates [68]. From this and other legislation, two key definitions of discrimination have emerged in AI literature: disparate treatment, which refers to intentionally discriminatory practices targeted at individuals based on their protected attributes; and disparate impact, which encompasses practices that result in a disproportionately adverse effect on protected groups [67, 69, 70]. Although it might seem straightforward to meet these standards by simply removing protected attributes from the training data, many situations still arise where, despite these efforts—and sometimes because of them—unfair decisions are made. For instance, in natural language processing (NLP), a model might infer a candidate’s gender from their name, a phenomenon known as indirect discrimination or discrimination through features implicit in the dataset [69]. Therefore, to achieve fairness in a decision-making system, it is essential to address the underlying causes of bias in the data. This is discussed next.
II-C Causes of Bias
There are several common ways in which human bias can be transferred to an AI model [66, 1, 71]:
  1. 1.
    Training data: If the training data is biased, an AI system trained on it will learn and propagate that bias. This can occur through a skewed sample, where one group is disproportionately represented in achieving a particular outcome compared to another, or through tainted label outcomes (e.g., if the dataset was manually labeled by humans, any inherent human bias may be transferred to the labels).
  2. 2.
    Label definitions: The target label in AI represents the outcome a model aims to predict. Selecting a holistic or vague definition of the target outcome can result in a larger disparate impact. For example, if a manager is building a classification model to predict a job candidate as a “good” hire, various definitions could be applied to determine a “good” hiring decision [66]. Factors such as employee motivation, person-job fit (i.e., the alignment of an applicant to a job), predicted tenure, or person-environment fit (i.e., the compatibility of an applicant with the organization’s culture and workplace) typically determine how well a candidate fits an organization and their expected performance once hired [72, 73, 66]. However, selecting a label with a systematically disproportionate value for members of a protected class will bias the model’s predictions [66]. For instance, using predicted tenure as a target label, which may seem reasonable to an employer, might result in bias against women if historical data shows lower tenure rates for women compared to men [66].
  3. 3.
    Feature selection: Features are attributes that describe or contribute to an outcome that an AI model uses during prediction. While feature selection can improve model accuracy by reducing data complexity and learning time [74], it can also lead to bias if the statistical variation between individuals in protected groups is not well-represented for a particular feature [66]. Features may lack the necessary granularity for an AI model to learn accurate classification for protected groups [66]. For instance, academic credentials and the college from which a job candidate graduated are commonly used in hiring; however, if members of a protected group graduate at a disproportionately higher or lower rate from certain colleges, an AI model might incorrectly classify job candidates from the protected group, even if they are qualified for the job [66].
  4. 4.
    Proxies: Even with protected attributes removed from the training data, group membership can be inferred from other attributes, potentially leading to biased decisions. For instance, Amazon’s hiring application was biased even without using the gender attribute because the model could infer gender from the educational institution listed on the resumes of applicants (e.g., all-female or all-male colleges) [13].
  5. 5.
    Masking: If model developers or employers hold prejudicial views, they might collect data samples or engineer features that intentionally bias the AI model [66]. This practice is known as masking, where a data representation (e.g., using only coarse-grained model features that obscure fine-grained group details) is employed to justify AI model bias toward certain protected groups [66].
III Bias in AI-Based Recruitment
HR and I-O psychology literature typically divides the recruitment pipeline of an organization into four stages, as depicted in Figure 2 [21, 38, 75]. This section provides an overview of AI-driven practices, potential biases, and efforts to achieve fairness at each stage of the recruitment process.
Figure 2: Overview of steps in a typical recruitment pipeline: (1) candidate sourcing (after job analysis), (2) candidate screening, (3) candidate interviews, and (4) selection and offer, including post-decision evaluation.
III-A Sourcing
The first stage of the recruitment process begins with a job analysis, where an employer identifies a position within the organization and its required experience, knowledge, skills, abilities, and other pertinent characteristics (KSAOs). This stage also involves evaluating criteria for person-environment fit [72, 76, 77], which includes assessing how well an individual’s values and traits align with the company’s work culture.
Following the job analysis, the employer moves to sourcing candidates, which involves creating a pool of potential candidates by attracting job seekers to apply. This is accomplished by crafting job ads that reflect the identified KSAOs, conducting recruiter outreach, and leveraging employee referrals [38, 75].
In this context, AI-based methods have been developed for both job analysis and candidate sourcing stages. AI-based candidate sourcing platforms, such as HireEZ [78] and Arya [79], utilize AI to reach a broader and more diverse range of candidates [80]. Additionally, LLMs are employed for various tasks, including crafting job ads [81], conducting job analysis, and sourcing candidates through prompting-based interfaces like TalentGPT [82, 83]. Other AI-powered services, such as Textio, BrightHire, and SeekOut [84, 82, 85, 86], are also available to streamline the candidate sourcing process.
However, two prominent areas of bias can occur in AI-based sourcing software. The first area of bias is in the text of job ads, which have been found to use gendered keywords and phrases, as well as perpetuate age and racial biases [81, 87, 88, 89, 86]. To address this issue, services like Textio use AI to provide neutral language and reduce potential bias [81]. Unfortunately, many job ads are not tested for bias in their wording, which can negatively impact not only the time and resources spent by organizations to find qualified candidates but also the trust and self-efficacy, or confidence, of the jobseeker [90, 6, 4]. Biased job ads can also influence AI-driven job-ad recommendations, leading to skewed ads being shown to users, as observed in previous studies on Facebook and Google job ad recommendations [14, 12].
Job recommendation systems, used by jobseekers to browse job ads, heavily rely on AI to match resumes with ads in their database, often leading to biased recommendations. Examples of such platforms include LinkedIn [91] and Indeed [92], where job postings are aggregated for jobseekers to search, while recruiters frequently search for passive candidates—individuals who are not actively seeking a job but fit a job ad well [38]. The recommendation algorithms on these sites typically use collaborative filtering [93], content-based filtering [94, 93], or multi-objective ranking approaches [95]. However, job ad content can lead to biased recommendations, as an estimated 60% - 70% of job ads posted online disproportionately represent industries for high-skilled workers and STEM occupations [96]. Moreover, ad delivery systems can be skewed along gender and racial lines, as evidenced by Facebook’s job recommendation engine in 2019 [14]. Additionally, gender inequality was observed on job boards such as Indeed, Monster, and CareerBuilder in 2018, with 12 of the 35 studied job titles showing significant group unfairness [97].
To address such biases, various approaches have been developed. For example, Imana et al. distinguish between skew in ad delivery caused by protected attributes and skew due to differences in candidate qualifications [98]. Similarly, Chen et al. [99] tackle recency bias in AI-driven information retrieval, where recent content disproportionately influences model training, leading to bias. Additionally, Rus et al. [100] use a generative adversarial network to eliminate gender biases from word vector representations in job ad text and resumes. Meanwhile, Li et al. model the job recommendation task as a resource allocation problem, promoting fairness by maintaining consistency across users from different protected groups [101]. Despite these efforts to mitigate cognitive biases in job and candidate recommendations, such biases may persist through the training data [75].
III-B Candidate Screening
The next stage in recruitment, candidate screening (or candidate matching), involves filtering an initial pool of candidate profiles, such as resumes, based on person-job fit. During this process, large companies that handle thousands of applications often use ranking and screening tools. Examples of AI-based screening methods include those provided by HireVue [102], Plum [103], and SeekOut [82]. These services offer resume ranking and filtering systems, assessments of candidate job experience and knowledge, as well as personality evaluations [77, 38, 104, 72].
However, similar to other areas of AI use, screening models are susceptible to bias, as observed in past applications such as Amazon’s resume ranking application [13, 105]. Institutional and systemic biases are often reflected in resume screening and assessment models because the training data is based on previous hiring decisions, which may lack representation from different protected groups [75]. Additionally, these systems, often based on NLP, may generate harmful language, perpetuate stereotypes, or create unequal representation of different groups [106]. For instance, in 2016, popular word embeddings, which are frameworks used to represent text data as vectors for training in neural networks, were found to exhibit gender stereotypes due to training on Google News articles [106]. Later, GPT-3 was found to generate strong negative stereotypes against religious groups [34], and other LLMs like GPT-4 and BERT have also come under scrutiny for bias [107]. These biases can negatively impact both jobseekers and organizations. If less-qualified candidates are hired due to bias, it can lead to additional time and resources being required for training and onboarding, potentially resulting in lower employee retention and job satisfaction rates [90].
Recognizing the importance of screening processes, many researchers have focused on developing methods to detect and mitigate biases. For example, Pena et al. [108, 109] propose a testbed for resume screening algorithms to identify bias in candidate scoring across five scenarios with different protected attributes. Similarly, Deshpande et al. [110] tackle bias by reweighting tf-idf values based on fairness metrics when matching resumes with job descriptions. Burke et al. [111] introduce a mitigation method through candidate ranking. While these techniques represent an initial effort in bias mitigation, non-textual features within screening models, such as images (commonly found in online profiles), and key non-protected attributes like educational background may still contribute to perpetuating bias. This issue, where non-textual features such as visual and auditory cues contribute to bias, is particularly prevalent in the screening of candidate video or phone interviews.
III-C Interviewing Candidates
The third stage in the recruitment process involves interviewing candidates to assess their KSAOs, work values, personality, and other traits in detail. Over the past decade, the recruitment process has increasingly relied on AI to evaluate these traits during interviews. Various AI modalities have been utilized, including chatbot systems for candidate interaction, automated video and phone interviews, and online question-answer assessments [102, 23, 32, 75, 23]. For instance, platforms like HireVue [102] and BrightHire [85] heavily depend on AI to evaluate video interviews, scoring candidates based on personality traits and their fit for the job.
However, in the realm of video- and image-driven AI applications, concerns have arisen regarding potential biases, as highlighted by various studies [112, 113, 114]. This issue became prominent in 2021 when HireVue ceased the use of its video-based candidate scoring systems, citing the absence of transparency in AI decision-making as a primary reason [32]. Additionally, in 2020, Illinois enacted the Artificial Intelligence Video Interview Act, the first regulation addressing AI in the interviewing of job candidates [115]. This law mandates that employers must inform applicants that their interview will be scored by an AI-driven system, provide the characteristics used by the AI system in scoring, and obtain consent from each applicant before the interview [115]. Furthermore, research into AI-based video interviewing systems and chatbots has revealed biases originating from the training data [116, 117]. For instance, investigations into the First Impressions dataset, used in the 2017 ChaLearn challenges at CVPR, uncovered biased pairwise labels assigned to participants’ personality traits based on gender and race [118, 119]. When these labels were used to construct a model, the biases were further amplified through the model’s predictions [118, 119].
To promote fair candidate screening, researchers have implemented various strategies to address these biases. For example, Yan et al. have proposed an approach using adversarial learning to mitigate bias according to fairness metrics in personality trait prediction [119]. Additionally, Booth et al. [120] have studied bias measurement in affective computing problems, particularly in the context of mock video interviews. Significant efforts have also been made to develop chatbot interviewing systems that ensure fairness. Researchers have explored the use of fairness-aware classifiers to mitigate harmful text content and identify biased prompts from users [121]. The survey by Xue et al. [121] provides a comprehensive overview of these approaches and potential biases in chatbot systems.
III-D Selection and Evaluation
The final stage of the recruitment process involves selecting one or more job candidates, followed by negotiating an offer, which is then used to evaluate the organization’s overall recruitment process. AI has been incorporated in several ways to assist not only in selection but also in the negotiation of job offers. For example, AI is used during job offer negotiations to provide salary and training recommendations, thereby aiding in the onboarding process of newly hired employees [75]. This is exemplified by software like Oracle Recruiting Cloud, which not only forecasts the probability of a job candidate accepting an offer but also allows employers to experiment with various salary structures, benefits packages, and other incentives to enhance the likelihood of acceptance [75]. Additionally, Pactum AI [122] automates various negotiation processes, such as salary negotiations and contract discussions, leveraging a multitude of factors, including job market outlook.
Although such AI seeks to create a fairer negotiation process by eliminating human biases, other forms of bias can still occur. For instance, a 2019 UpTurn report highlighted that Oracle and other AI-driven negotiation tools might reinforce asymmetric information between the job candidate and employer, potentially leading to an unfair process for the candidate [75]. Additionally, these tools may conflict with legislation prohibiting the use of salary history in hiring decisions and the Equal Pay Act of 1963, which mandates equal pay for individuals in the same position [75, 62].
Job candidates who are not hired often face a lack of transparency regarding the AI-driven decision-making processes of employers. According to a survey conducted by The Talent Board, 53.5% of candidates reported receiving no feedback after the screening and interview stages, while 69.7% received no feedback after being turned down [123]. Although employers are not legally obligated to provide explanations [123], the absence of feedback can negatively impact the perceived fairness and trustworthiness of an AI-based recruitment process. Transparency is crucial at every stage of the process, including communicating the job outcome, as it enables candidates to learn and improve for future interviews. Previous studies have demonstrated that providing explanations for organizational decisions can significantly enhance trust [41, 38]. Recently, many approaches, as further described in Section IV-B, have prioritized model interpretability and explainability to enhance transparency in AI-based decision-making, a principle that can be extended to AI-driven hiring decisions.
After finalizing hiring decisions, organizations evaluate their recruitment process based on three key criteria: validity, utility, and fairness. Validity ensures that assessments and interviews accurately reflect job KSAOs and predict job performance [38]. For instance, organizations often establish a criterion-referenced cut score as a cut-off for employee selection, based on the desired level of performance for a new employee [38]. Compliance with government guidelines, as described in Section II-B, is also essential to validity [38]. Next, utility assesses the cost-benefit ratio of staffing, which can be influenced by earlier recruitment stages such as employee outreach and job marketing. Finally, the fairness criteria is evaluated using employee selection guidelines described in Section II-A. However, as AI becomes more integrated into each stage of the process, it is necessary to incorporate bias detection and mitigation methods where guidelines may not adequately address algorithmic bias. For example, the uniform guidelines on employee selection could be substantially affected by the sample size of different groups. Work by Peng et al. [124] demonstrates how AI-human, or hybrid, decisions where AI assists humans, can still carry bias. Moreover, legislative definitions of fairness and employee selection guidelines may not be sufficiently detailed or up-to-date to address recent technological advancements.
IV Bias Detection and Mitigation
To ensure fairness in AI-driven hiring, it is crucial to implement methods for detecting and measuring bias and mitigating it at each stage of the hiring pipeline. In this section, we review the existing work in AI fairness, focusing on methods for measuring bias, strategies for mitigating bias, and the current limitations and challenges in this field.
IV-A Fairness Metrics
In AI fairness research, various definitions of fairness have been proposed to assess AI systems effectively. In this subsection, we explore core fairness metrics particularly relevant to recruitment, as defined in prior research [66, 71, 74].
  1. 1.
    Fairness through unawareness: This metric serves as a baseline by stating that an algorithm is fair if it does not explicitly use protected attributes $A$ in its training [125]. The classifier $C$, defined as $C:X\setminus A\to Y$, operates on a set of features $X$ excluding any protected attributes $A$. Here, $Y$ represents the potential outcome classes. While this approach aims to prevent direct discrimination by ensuring the classifier is ’unaware’ of protected characteristics, it has notable shortcomings. In the hiring context, where $A$ might vary significantly between different occupations, expert knowledge is essential to accurately identify and exclude these attributes. Moreover, even when $A$ is omitted from $X$, other features in $X$ might act as proxies to $A$, leading to a classifier that is nominally fair while still harboring discriminatory information. For example, while gender may not be explicitly included, participation in a women’s coding bootcamp listed in job history could still suggest the candidate’s gender, thus acting as a proxy and influencing the hiring decision. Additionally, educational background might act as a proxy for socioeconomic status or ethnic background, especially if certain schools predominantly serve specific demographic groups.
  2. 2.
    Demographic parity: Also known as statistical parity, this metric stipulates that the acceptance rates for different demographic groups should be approximately equal. This requirement aligns with fair hiring practices mandated by laws such as the four-fifths rule (see Section II-B). Mathematically, let $C$ be a classifier predicting a binary outcome $Y$. For any two groups $a$ and $b$, the acceptance rates must be within a specified percentage $p$ of each other, where $\epsilon=\frac{p}{100}$ and $\epsilon\in[0,1]$ [71]:
    $\left|P_{a}\{C=1\}-P_{b}\{C=1\}\right|\leq\epsilon.$ (1)
  3. 3.
    Accuracy parity: Also known as equality of opportunity, this metric ensures that the true positive rate (the proportion of actual positives correctly identified by the model) is equal across different demographic groups [70]. This form of fairness seeks to equalize the likelihood of a qualified candidate from any group being hired (i.e., $C=1$ when $Y=1$). This concept is mathematically represented by the equation [1]:
    $P_{a}\{C=1\mid Y=1\}=P_{b}\{C=1\mid Y=1\}.$ (2)
    This metric does not address the trade-off between false positives and false negatives directly but focuses on ensuring that all candidates who are qualified have an equal chance of being selected, regardless of group membership.
  4. 4.
    Predictive rate parity: Also known as positive predictive parity and negative predictive parity, this metric ensures that the credentials of a candidate are consistent with the model’s predictions across different groups [1]. Predictive rate parity is satisfied if both positive and negative outcomes are equally well predicted across groups. Positive predictive parity requires that the probability of a positive classification given a positive label is the same for all groups:
    $P_{a}\{C=1\mid Y=1\}=P_{b}\{C=1\mid Y=1\}.$ (3)
    Negative predictive parity requires that the probability of a negative classification given a negative label is the same across groups:
    $P_{a}\{C=0\mid Y=0\}=P_{b}\{C=0\mid Y=0\}.$ (4)
    These conditions ensure that the model’s predictions do not disproportionately benefit or penalize any group, regardless of the actual outcome.
  5. 5.
    Counterfactual fairness: Motivated by the idea that decisions made by an algorithm should not change if a protected attribute $A$ were hypothetically altered, counterfactual fairness enhances transparency by demanding robustness in predictions against changes in $A$. Defined by Kusner et al. [125], a classifier $C$ achieves counterfactual fairness if, for any individual with features $X$ including a protected attribute $A$, the outcome remains the same in both the actual and a counterfactual scenario where $A$ is altered. The mathematical definition is:
    $P\{C_{A\leftarrow a}(X)=y\mid X,A=a\}\\ =P\{C_{A\leftarrow a^{\prime}}(X)=y\mid X,A=a\},$ (5)
    where $a$ represents the actual value and $a^{\prime}$ represents a counterfactual value of $A$, and $y$ is the potential outcome. This ensures the prediction is independent of $A$, except for its legitimate influences via other non-protected attributes.
  6. 6.
    Individual fairness (versus group fairness): Unlike group fairness metrics such as demographic parity, accuracy parity, and predictive rate parity which ensure fairness by equalizing statistical measures across groups, individual fairness focuses on fairness at the individual level. This metric asserts that similar individuals should receive similar outcomes. It addresses the key limitation of group fairness metrics, which ensure fairness across groups but may still permit unfairness within groups [104]. Specifically, individual fairness ensures that any two individuals $i$ and $j$ who are similar with respect to the task at hand should be treated similarly [126].
  7. 7.
    Preference-informed individual fairness: This fairness metric extends individual fairness by considering the diverse preferences of individuals, such as job location, industry, salary range, and work hours. Preference-informed individual fairness (PIIF) integrates the concept of envy-freeness, which ensures that no individual would prefer another’s outcome to their own under similar conditions [126]. Thus, PIIF demands that outcomes not only meet individual fairness criteria but also align with the individual preferences to mitigate envy, ensuring each decision made by the model is both fair and personally satisfactory to the individuals involved.
  8. 8.
    Multi-sided fairness: Modern recruitment platforms necessitate a multi-sided approach to fairness, considering the provider (i.e., employers posting job advertisements), the consumer (i.e., jobseekers), and the platform itself (i.e., the job posting recommendation or screening system) [127]. Prioritizing the interests of only one party, such as focusing solely on the employer (producer-focused) or the jobseeker (consumer-focused), can lead to unfair outcomes for the other. For instance, a consumer-focused approach, or C-fairness, might mandate that job ads display similar salary ranges to all users regardless of protected group membership. Conversely, a producer-focused approach, or P-fairness, might require that ads from minority-owned businesses be shown to qualified jobseekers at comparable rates to those from other businesses [128, 129]. This concept has been applied in various online platforms such as e-commerce and ride-hailing, where fairness is treated as a multi-objective optimization problem, balancing accuracy and fairness [127, 130, 128, 129].
  9. 9.
    Procedural fairness: Procedural fairness, one of the three types of organizational justice discussed in Section II-A—alongside distributive and interactional justice—focuses on the fairness of processes within an organization, such as recruitment. Grgic-Hlaca et al. [44] propose using procedural fairness to evaluate recruitment processes. Specifically, the fairness of a classifier $C$ and its feature set $X$ is assessed based on the proportion of users who perceive the use of certain features in the decision-making process as fair. Essentially, this approach allows users to actively evaluate the fairness of decisions. However, perceptions of fairness may vary over time and are influenced by the users’ awareness of potential disparities caused by specific features in the model [44].
  10. 10.
    Fairness in regression: While many established fairness metrics focus on classification problems where decisions are typically binary (e.g., accept or reject), regression models are also prevalent in hiring processes, such as scoring video interviews [131]. These models, however, have received less attention regarding fairness. According to Barocas et al. [70], fairness in regression is guided by three principles:
    • Independence — ensuring that a sensitive attribute $A$ and the regression output $R$ are independent, denoted as $R\perp A$, implying that the sensitive attribute has no influence on the model’s output.
    • Separation — allowing for some correlation between $A$ and the target variable $Y$ if it is justifiable, denoted as $R\perp A\mid Y$.
    • Sufficiency — ensuring that the regression output adequately captures the influence of $A$ on $Y$, aiming for parity among individuals based on their relevant characteristics, denoted as $Y\perp A\mid R$ [66].
    Additionally, fairness can be quantified using statistical measures such as the Kullback-Leibler divergence to compare predictions across different groups [30].
Each of the above metrics addresses different aspects of fairness and is applicable in various scenarios within the recruitment process. Their effective implementation can help mitigate biases and ensure a more equitable hiring process.
IV-B Bias Mitigation
After detecting bias within an AI system, the next step involves implementing strategies to mitigate these biases effectively. This subsection outlines three primary categories of bias mitigation techniques as identified in existing literature [132]:
  1. 1.
    Pre-processing: This approach involves modifying the dataset before the model training process begins. Techniques in this category, such as reweighting and optimized pre-processing, adjust features and labels in the dataset to align with fairness criteria before the model is trained and used for prediction [132]. These methods aim to remove or modify bias-inducing factors in the data, such as unrelated protected attributes. Toolkits like AI Fairness 360 offer a range of pre-built functions for these pre-processing tasks [132]. However, while effective at data-level adjustments, pre-processing does not rectify biases inherent to the core modeling process, which may still require attention if biases persist after dataset adjustments.
  2. 2.
    In-processing/optimization: Unlike pre-processing, in-processing methods involve optimizing the model during training to meet specific fairness definitions through constraints or modifications in the model’s design. This could, for instance, involve implementing accuracy parity by ensuring that a proportionate number of qualified candidates from different demographic groups are hired. In-processing techniques fall into two categories: explicit methods, which directly alter the training objectives, and implicit methods, which modify the representation of data within the model [133]. A comprehensive discussion on these methods can be found in the survey by Wan et al. [133]. The primary advantage of in-processing is its effectiveness in optimizing fairness metrics; however, this may sometimes come at the expense of reduced model accuracy [134]. Notably, in-processing presents challenges such as the necessity to alter the model architecture, which may not always be feasible, particularly in outsourced recruitment scenarios. Additionally, if the dataset’s target labels are inherently biased due to systemic issues, these biases may continue to be propagated, despite the application of fairness constraints during model training.
  3. 3.
    Post-processing/transparency: Post-processing techniques focus on adjusting the outputs of an already-trained model to satisfy fairness criteria. For instance, this might involve modifying the classification thresholds used to make hiring decisions. The primary advantage of this approach is that it does not necessitate changes to the underlying model architecture. Additionally, these methods often incorporate elements of algorithmic transparency, such as providing explanations for decisions through counterfactual reasoning or analysis of feature importance. Revealing the decision-making process enables a thorough assessment of its fairness and allows for necessary adjustments if biases are identified. Moreover, providing clear explanations helps job candidates understand potential discrepancies or deficiencies in their applications, which can be beneficial for their job-seeking future endeavors. Research has shown that transparent feedback can significantly improve candidates’ perceptions of fairness within the recruitment process [135].
Numerous ML toolkits have been developed to facilitate the integration of fairness algorithms at various stages of the ML pipeline. These toolkits, comprehensively reviewed in the survey by Lee et al. [136], support the implementation of pre-processing, in-processing, and post-processing methods. Their availability enables organizations to more easily embed fairness into their AI-driven hiring processes, thereby helping ensure that their recruitment practices meet established fairness standards.
IV-C Observations, Limitations, and Future Challenges
Here, we outline four key limitations in AI-driven hiring that underscore the need for ongoing research to enhance the fairness of AI in recruitment processes:
  1. 1.
    Selecting metrics and auditing practices: As organizations implement fairness practices, a pivotal challenge arises in selecting appropriate fairness metrics, which can vary significantly across different applications. In recruitment, for example, it is crucial to consider both candidate qualifications and the representation of protected groups. Additionally, the difficulty of meeting multiple fairness criteria simultaneously [134] suggests that relying on a single metric might ignore other significant disparities. Equally concerning is the absence of standardized benchmarking and comprehensive auditing procedures in current legislation, which are vital for consistently ensuring fairness in the use of AI tools in hiring [109].
  2. 2.
    Job-specific requirements and long-term fairness: The criteria for measuring fairness must adapt to the unique requirements of different jobs. For example, while religion is generally a protected attribute, it may be relevant for leadership roles within religious institutions. This underscores the challenge of using a single model to predict protected attributes universally across all job types, especially where specific attributes are crucial. Additionally, the need for fairness extends to job advertisements and their recommendations, which must align with the detailed requirements specified in the job descriptions. As the job market evolves, the models and auditing mechanisms must also be dynamic, adapting to long-term changes in the market. This dynamic aspect of fairness, focusing on long-term fairness impacts and adaptability, has been underexplored in the fairness literature, which has traditionally focused on static fairness assessments.
  3. 3.
    Auditing large foundation models: Large foundation models like ChatGPT, Whisper, and DALL-E are increasingly employed across various hiring tasks due to their robust performance in diverse scenarios. These models utilize extensive training datasets comprising text, images, and audio. However, their widespread use raises significant ethical concerns due to the current lack of comprehensive regulations that ensure fairness throughout the model training and deployment phases. Key issues include preventing the leakage of sensitive data when generative models create content and promoting diversity in training datasets to minimize the perpetuation of stereotypes and harmful biases, thereby safeguarding equal treatment of all individuals [137].
  4. 4.
    Hiring decision transparency: Many applicants never receive feedback if screened out by an AI-based system, and those who do often get responses lacking detailed justification. This lack of detailed feedback on unmet requirements or the features influencing the AI’s decision undermines interactional justice and erodes trust. Enhancing the interpretability and explainability of ’black-box’ AI models is crucial for fostering transparency and fairness in AI-driven hiring. Despite emerging legislation aimed at improving transparency in AI-driven decisions, such initiatives are still limited and often lack robust enforcement mechanisms.
These outlined challenges highlight the complex, evolving landscape of AI-driven hiring, underscoring the critical need for continued innovation and rigorous evaluation to advance fairness and transparency in recruitment practices.
V Conclusion
In this paper, we have systematically reviewed the existing literature on fairness in AI-based recruitment systems, elucidating the complexities and challenges at the intersection of AI research and development, I-O psychology, and legislative frameworks. We began by defining essential concepts such as fairness, trust, and justice, grounding our discussion in relevant literature to provide a foundational understanding for analyzing bias. We then detailed the recruitment pipeline, illustrating the application of AI at each stage to both detect and mitigate bias, and highlighted the current methodologies and their limitations as outlined in the fairness literature, specifically tailored to AI-driven recruitment.
We identified significant areas in need of future research, including the development of models that cater to job-specific requirements, the creation of fair job advertisements to engage a broader audience, and enhancing transparency in recruitment decisions. Additionally, pinpointing the origins of bias remains a pivotal challenge for the advancement of fair AI practices in recruitment.
The prevalence of bias in AI recruitment models poses considerable risks to organizations, job seekers, and the broader job market. Despite the enhanced efficiency provided by AI in screening vast numbers of applications, the propensity of these systems to perpetuate existing human biases is a significant concern. Addressing these issues requires robust enhancements in AI recruitment practices, incorporating thorough fairness auditing, transparent AI decision-making processes, and measures ensuring that job-specific requirements do not introduce further biases. Moreover, as the development of large foundation models like ChatGPT, Whisper, and DALL-E continues, the protection of applicant data privacy becomes increasingly imperative during model training and refinement phases.
Despite these challenges, the advancements in large foundation models offer unprecedented opportunities to transform recruitment processes worldwide. This potential drives the need for concerted efforts among organizations, researchers, and policymakers to ensure that the evolution of the recruitment landscape is both innovative and equitable, potentially leading to more dynamic workplaces and a more engaged workforce.
Acknowledgment
This material is based upon work supported by the National Science Foundation under Grant No. 1936857.